OpenStudy (anonymous):

Please help!!! What is equal to (x+9)(x-2)/(x-2)(x+2) a. x-2/2+x b.x+9/x-2....my answer c.x+9/2+x d.2+x/x+9

OpenStudy (jhannybean):

\[\frac{(x+9)\color{red}{(x-2)}}{(x+2)\color{red}{(x-2)}}\] Treat the highlighted ortion as 1. \[\frac{(x-2)}{(x-2)} = 1\]

OpenStudy (jhannybean):

Substitute 1 in place of that fraction, what will you get?

OpenStudy (anonymous):

1?

OpenStudy (jhannybean):

What I meant is\[\frac{(x+9)\color{red}{(x-2)}}{(x+2)\color{red}{(x-2)}} \implies \frac{(x+9)}{(x+2)}\color{red}{(1)} = ~?\]